# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 10 Perspectives
This review centers on the Mpemba effect, beginning with the original research conducted by Mpemba and Osborne in 1969 [1] and providing a historical context that traces back to Aristotle. Our focus, however, is on the significant advances made in this field over the past decade. After discussing the various challenges and efforts made to explain the Mpemba effect in water—highlighting its unique properties in Sec. 1.1—we examine the range of observations and characterizations of the Mpemba effect and closely related phenomena. The different types of Mpemba effects defined in Sec. 2 that can be understood within Markovian frameworks are crucial for understanding the effect. They have sparked considerable interest in identifying similar anomalous-relaxation phenomena across various systems, from single-particle interactions to many-body physics, in both classical and quantum settings.
@@
The question of how different effects, frameworks, and experiments are related poses a significant challenge for the coming years. For instance, can the Markovian framework discussed in Sec. 2 explain the phase transition effects covered in Sec. 3? Additionally, can the intuitive single-particle perspective from experiments mentioned in Sec.  6 be applied to macroscopic systems such as water, magnetic alloys, and others? Furthermore, is it possible to extend the Boltzmann-Fokker-Planck equation and the framework outlined in Sec. 4 to make predictions about macroscopic materials such as water? It is our hope that a framework encompassing all these phenomena can be found, one that goes all the way from making ice cream in Tanzania to preparing states for quantum computing, and explains why—sometimes—the longer path can nevertheless be faster.
@@ 11 Acknowledgements
MV would like to express gratitude to Massimo Vergassola for advising that it is time to write a review of the Mpemba effect. We greatly appreciate the contributions of our coauthors and the vibrant scientific community surrounding the Mpemba effect. Special thanks to Siddharth Sane for generating Fig. 7. This material is based on work supported by the National Science Foundation under Grant No. DMR-1944539. MV would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme “Stochastic systems for anomalous diffusion”, where part of the work on this paper was undertaken. This work was supported by EPSRC grant no EP/R014604/1. JB was supported by Discovery Grants from the National Science and Engineering Research Council (NSERC, Canada). OR acknowledges financial support from the ISF, grant no. 232/23, and by the Minerva foundation with funding from the Federal German Ministry for Education and Research. AL was supported by the Project PID2021-128970OA-I00 financed by MICIN/AEI/10.13039/501100011033 and FEDER a way to make Europe and by C-EXP-251-UGR23 European Regional Development Fund, Junta de Andalucía-Consejería de Economía y Conocimiento.

### v3

@@ 10 Perspectives
This review centers on the Mpemba effect, beginning with the original research conducted by Mpemba and Osborne in 1969 [1] and providing a historical context that traces back to Aristotle. Our focus, however, is on the significant advances made in this field over the past decade. After discussing the various challenges and efforts made to explain the Mpemba effect in water, highlighting its unique properties in Sec. 1.1, we examine the range of observations and characterizations of the Mpemba effect and closely related phenomena. The different types of Mpemba effects defined in Sec. 2 that can be understood within Markovian frameworks are crucial for understanding the effect. They have sparked considerable interest in identifying similar anomalous-relaxation phenomena across various systems, from single-particle interactions to many-body physics, in both classical and quantum settings.
@@
The question of how different effects, frameworks, and experiments are related poses a significant challenge for the coming years. For instance, can the Markovian framework discussed in Sec. 2 explain the phase transition effects covered in Sec. 3? Additionally, can the intuitive single-particle perspective from experiments mentioned in Sec. 6 be applied to macroscopic systems such as water, magnetic alloys, and others? Furthermore, is it possible to extend the Boltzmann-Fokker-Planck equation and the framework outlined in Sec. 4 to make predictions about macroscopic materials such as water? It is our hope that a framework encompassing all these phenomena can be found, one that goes all the way from making ice cream in Tanzania to preparing states for quantum computing, and explains why—sometimes—the longer path can nevertheless be faster.
@@ 11 Acknowledgements
MV would like to express gratitude to Massimo Vergassola for advising that it is time to write a review of the Mpemba effect. We greatly appreciate the contributions of our coauthors and the vibrant scientific community surrounding the Mpemba effect. Special thanks to Siddharth Sane for generating Fig. 7. Also, we appreciate the critical reading and comments of the manuscript by Antonio Prados and especially Andrés Santos. This material is based on work supported by the National Science Foundation under Grant No. DMR-1944539. MV would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme “Stochastic systems for anomalous diffusion”, where part of the work on this paper was undertaken. This work was supported by EPSRC grant no EP/R014604/1. JB was supported by Discovery Grants from the National Science and Engineering Research Council (NSERC, Canada). OR acknowledges financial support from the ISF, grant no. 232/23, and by the Minerva foundation with funding from the Federal German Ministry for Education and Research. AL was supported by the Project PID2021-128970OA-I00 financed by MICIN/AEI/10.13039/501100011033 and FEDER a way to make Europe and by C-EXP-251-UGR23 European Regional Development Fund, Junta de Andalucía-Consejería de Economía y Conocimiento.